The compound V2Sx is 38.85%V by mass. What is x
1
Expert's answer
2011-05-05T12:46:50-0400
The molar weight of the compound is (51*2 + 32 + x) = 134+x
The part of V can be calculated as 51*2 / (134+x) = 0.3885, thus
x = 102/0.3885 - 134 =& 128.54.
The molecular weight of X is approximately corresponds to Tellurium